Zoe needs 30,000 for college in 12 years. Will she have enough money for college if she invests 7,250 at 10% intrest, compounded annually, now?

Answer by stanbon(75887)

Zoe needs 30,000 for college in 12 years. Will she have enough money for college if she invests 7,250 at 10% intrest, compounded annually, now?
-------
A(t) = P(1+(r/n))^(nt)
----
A(12) = 7250(1+(0.1/1))^(1*12)
----
= 7250(1.1)^12
---
= 7250*3.1384
---------
= $22,753.61
------------------
Not going to have enough
=====================
cheers,
Stan H.
===================